OpenStudy (anonymous):

Solve by factoring: 2x^2-5x-3=0?

OpenStudy (aroub):

(2x-3)(x-1)

OpenStudy (anonymous):

2x^2-6x+x-3 = 0 2x(x-3)+1(x-3) = 0 (2x+1)(x-3) = 0 x=-1/2,x=3

OpenStudy (anonymous):

which one is correct?

OpenStudy (anonymous):

aroub can u explain this ?

OpenStudy (aroub):

lol idk!

OpenStudy (aroub):

yes sure!

OpenStudy (anonymous):

thanks

OpenStudy (aroub):

ok , ill explain my answer

OpenStudy (aroub):

but what do u want to know exactly?

OpenStudy (anonymous):

how you got the factors

OpenStudy (aroub):

ok , there are many different ways to do it!!!! but ill tell u my way

OpenStudy (aroub):

ok ill use ur example 2x^2-5x-3 ok u look at 3 and u say to urself what times what gives u 3, the only numbers u have is 3*1, then u look at 2 what times what gives u 2x^2 its 2x and x so u put them in brackets ex. (2x )(x ) then u look at the signs here in the example u have negative and negative the first sign u see it u put it in the first bracket (2x- )(x ) the second/ lol never mind ill write it on a paper and then put it in here for u to see it okay can u wait just another 5 min please i really want to help u !!

OpenStudy (anonymous):

thanks again if u dont mind please post it i ll never forget your help aroub thanks a 100 times

OpenStudy (aroub):

anytime my freind u seem like a very good , nice , kind person! ill sure like to help these kind of people!!!

OpenStudy (aroub):

but i want to post other example for u to understand more... okay?

OpenStudy (aroub):

oh and tell me if u can see the numbers ... lol

OpenStudy (anonymous):

when u r getting time please post it,i am going once again thanks

OpenStudy (aroub):

okay inshallah we u come back u will see it posted okay , byyeee!!! ::) :)

OpenStudy (aroub):

when**

OpenStudy (anonymous):

inshallah tomorow

OpenStudy (aroub):

wait can u wait 2 more min pleasee

OpenStudy (anonymous):

okay :)

OpenStudy (aroub):

LOL INSHALLAH TOMORROW TOMORROW bec. the numbers are not really showing ill fix it and post it on here

OpenStudy (aroub):

SORRYY!! :( :(

OpenStudy (aroub):

oh sry but r u muslim??

OpenStudy (anonymous):

my name is jumaliya can u guess

OpenStudy (aroub):

then u are!?1!!!!

OpenStudy (anonymous):

yea

OpenStudy (aroub):

i know jumana but not jumaliya

OpenStudy (aroub):

oh great! bec. i keep saying inshallah not realizing ALHAMDILLAH u are!!! :D

OpenStudy (anonymous):

bye arounb thank you very much for all the helps

OpenStudy (aroub):

OKAY BYE/ ANYTIME XX